Find the complex exponential Fourier series coefficients for:
a. x(t)=3cos(4w0t)
b. x(t)=sin^2(t)
2. Find the DTFT of:
x(n)={1 0<=n<=3
0 otherwise}
1
Expert's answer
2020-05-11T14:26:46-0400

a. Find the complex and exponential form:


"x(t)=3\text{cos}(4w_0t)\iff0+j\frac{3}{\sqrt{2}}\iff\frac{3}{\sqrt{2}}e^{j90^\circ}."

Fourier series coefficients from -π to π:


"a_0=\frac{\int^\pi_{-\pi}\text{cos}(4\omega_0t)dx}{\pi},\\
a_n=\frac{\int^\pi_{-\pi}\text{cos}(4\omega_0t)\text{cos}(tk)dx}{\pi},\\
b_n=\frac{\int^\pi_{-\pi}\text{sin}(tk)\text{cos}(4\omega_0t)dx}{\pi},\\
k=1,2,3,..."

b. Find the complex and exponential form:


"x(t)=\text{sin}^2(t)=\frac{1}{2}-\frac{\text{cos(2t)}}{2}\iff\frac{1}{2}-\frac{j}{2\sqrt2}\iff\\\iff\frac{1}{2}-\frac{e^{j90^\circ}}{2\sqrt2}."

Fourier series coefficients from -π to π:


"a_0=\frac{\int^\pi_{-\pi}\text{sin}^2(t)dx}{\pi},\\
a_n=\frac{\int^\pi_{-\pi}\text{sin}^2(t)\text{cos}(tk)dx}{\pi},\\
b_n=\frac{\int^\pi_{-\pi}\text{sin}^2(t)\text{sin}(tk)dx}{\pi},\\
k=1,2,3,..."

2. Find the DTFT of


"x(n)=\begin{Bmatrix}
 1,& \text{if }1\leq n\leq3 \\
 0&\text{otherwise}
\end{Bmatrix}"

"X(\Omega)=\sum^3_1e^{-j\Omega n}=e^{-3j\Omega}(e^{j\Omega}+e^{2j\Omega}+1)."